# Compatible $E$-differential forms on Lie algebroids over (pre-)multisymplectic manifolds

### v3

@@ 3.2 Gauge transformations
Recall that $\iota^{n+1}_{\rho}\omega$ is closed with respect to the Lie algebroid differential ${{}^{E}\mathrm{d}}$, ${{}^{E}\mathrm{d}}(\iota^{n+1}_{\rho}\omega)=0$ from consistency with Equation (11) and (12). Obvious equivalence relation of compatible $E$-$n$-forms is difference of ${}^{E}\mathrm{d}$-exact terms.
Definition 3.4
For a Lie algebroid $E$, two compatible $E$-$n$-forms $J$ and $J^{\prime}$ are equivalent if there exists an $E$-$(n-1)$-form $K\in\Gamma(\wedge^{n-1}E^{*})$ satisfying
$J^{\prime}=J+{}^{E}\mathrm{d}K.$ (14)
@@
${{}^{E}\mathrm{d}}(\iota_{\rho}^{n}\lambda)=\iota_{\rho}^{n+1}\mathrm{d}\lambda.$ (15)
@@
${{}^{E}\mathrm{d}}J^{\prime}$ $=$ $-\iota_{\rho}^{n+1}\omega^{\prime}=-\iota_{\rho}^{n+1}\omega-(\iota_{\rho})^{n+1}\mathrm{d}\lambda=-\iota_{\rho}^{n+1}\omega-{{}^{E}\mathrm{d}}(\iota_{\rho}^{n}\lambda)$ (16)
$=$ ${{}^{E}\mathrm{d}}(J-\iota_{\rho}^{n}\lambda).$
@@
In Section 4, we discuss gauge equivalence of the twisted Poisson structure as an example in Example 4.2.
@@ 4 Examples
The compatible $E$-$n$-form (11) is considered as a universal compatibility condition of a Lie algebroid structure with a pre-multisymplectic form. It appears in various geometries. In this section, we list up examples.
Example 4.1 (Liouville $1$-forms)
Take $n=1$ and let $(M,\omega)$ be a symplectic manifold and $E=TM$ be a tangent bundle. $TM$ is a tangent Lie algebroid with the anchor map $\rho=\mathrm{id}$ (Example 2.2). An $E$-$1$-form $J$ is a normal differential form on $M$, and a Lie algebroid differential ${}^{E}\mathrm{d}=\mathrm{d}_{\mathrm{dR}}$ is the de Rham differential. The compatible condition (11) is
$\mathrm{d}_{\mathrm{dR}}J=-\omega,$ (17)
@@
Example 4.2 (Poisson and twisted Poisson structures)
Let $(\pi,H)$ be a twisted Poisson structure on $M$ (Example 2.5). Note that it reduces to a Poisson structure if $H=0$. The cotangent bundle $T^{*}M$ has a Lie algebroid structure as explained in Examples 2.4 and 2.5. Formulas of the twisted Poisson structures and induced Lie algebroids are in [40, 32]. Equation (6) is equivalent to
${}^{E}\mathrm{d}\pi=-\langle\otimes^{3}\pi,\,H\rangle,$ (18)
where ${}^{E}\mathrm{d}$ is the Lie algebroid differential of this Lie algebroid $T^{*}M$. Thus, $J=\pi$ is a compatible $TM$-$2$-form on a pre-$2$-plectic manifold with a pre-$2$-plectic form $H$.
Next let us analyze gauge equivalence. A gauge transformation of the twisted Poisson structure is $H^{\prime}=H+\mathrm{d}B$ with a $2$-form $B$. Then, the compatible $E$-$2$-form with respect to $H^{\prime}$ is $\pi^{\prime}=\pi-\langle\otimes^{2}\pi,\,B\rangle$ from Definition 3.5.
@@
In the standard definition, Equation (6) gives the Lie algebroid structure on $T^{*}M$ and a compatible $E$-$n$-form simultaneously. If the gauge transformation $\pi^{\prime}=\pi(1+\pi B)^{-1}$ consistently deforms the Lie algebroid structure defined by $(\pi^{\prime},H)$ and a compatible $E$-$n$-form to $J^{\prime}=\pi^{\prime}$ simultaneously.
However, our gauge transformation $\pi^{\prime}=\pi-\langle\otimes^{2}\pi,\,B\rangle$ only deforms the compatible $E$-$n$-form to $J^{\prime}=\pi^{\prime}$ under the fixed Lie algebroid constructed from $(\pi,H)$. The reason is for generalizations to higher pre-$n$-plectic cases. In a pre-$n$-plectic case, the simultaneous deformation of the Lie algebroid and the compatible $E$-$n$-form like the twisted Poisson structure does not work. The twisted Poisson structure is a very special case and our gauge transformation is applicable to general $E$-$n$-form on a higher $n$-plectic manifold.
Example 4.3 (Twisted $R$-Poisson structures)
A twisted $R$-Poisson structure $(\pi,H,J)$ on a smooth manifold $M$ is defined as follows [12]. Since $J$ is denoted by $R$ in the paper [12], it is called the $R$-Poisson structure.
$M$ is a smooth manifold and $\pi\in\Gamma(\wedge^{2}TM)$ is a Poisson bivector field. $H$ is a closed $(n+1)$-form, and $J\in\Gamma(\wedge^{n}TM)$ is an $n$-multivector field. As explained in Example 2.4, the Poisson bivector field $\pi$ induces a Lie algebroid structure on $T^{*}M$, where $\rho=-\pi^{\sharp}$ and the Lie bracket is given by Eq. (5). Under this Lie algebroid structure, $(H,J)$ is the twisted $R$-Poisson structure if it satisfies
${}^{E}\mathrm{d}J=(-1)^{n}\langle\otimes^{n+1}\pi,\,H\rangle.$ (19)
@@
Example 4.4 (Momentum maps)
We assume a smooth action of a Lie group $G$ on a symplectic manifold $(M,\omega)$. Let $\mathfrak{g}$ be a Lie algebra for $G$. The Lie group action induces the Lie algebra action $\rho:\mathfrak{g}\rightarrow TM$. A map $\mu:M\rightarrow\mathfrak{g}^{*}$ is called a momentum map if it satisfies
$\mathrm{d}\mu(e)$ $=$ $-\iota_{\rho(e)}\omega,$ (20)
$\rho(e_{1})\langle\mu,\,e_{2}\rangle$ $=$ $\langle\mu,\,[e_{1},e_{2}]\rangle,$ (21)
for $e,e_{1},e_{2}\in\mathfrak{g}$. Here $\langle$-,\,$\rangle$ - , - ⟩ is the pairing of $\mathfrak{g}^{*}$ and $\mathfrak{g}$. As explained in Example 2.3, the Lie group action induces an action Lie algebroid structure on a trivial bundle $E=M\times\mathfrak{g}$. The anchor map is induced from the map of the Lie algebra action as $\rho:M\times\mathfrak{g}\rightarrow TM$. The momentum map is regarded as a section of the dual trivial bundle $\mu\in\Gamma(M\times\mathfrak{g}^{*})$. Under the condition (20), Equation (21) is changed to [5]
${}^{E}\mathrm{d}\mu(e_{1},e_{2})$ $=$ $-\iota_{\rho}^{2}\omega(e_{1},e_{2}),$ (22)
@@
Proposition 4.1
A momentum map $\mu:M\rightarrow\mathfrak{g}^{*}$ is a compatible $E$-$1$-form for the action Lie algebroid induced from a Lie group action on $M$.
Example 4.5 (Homotopy moment(um) maps)
[17] Let $(M,{\omega})$ be an $n$-plectic manifold. Assume an action of a Lie group $G$ on $M$ as in the case of the momentum map. The action of $G$ induces the corresponding infinitesimal Lie algebra action on $M$ as vector fields, $\rho:\mathfrak{g}\rightarrow TM$. We consider the space, $\Omega^{k}(M,\wedge^{n-k}\mathfrak{g}^{*})=\Omega^{k}(M)\otimes\wedge^{n-k}\mathfrak{g}^{*}$, and two differentials $\mathrm{d}:=\mathrm{d}\otimes 1$ and $\mathrm{d}_{\mathrm{CE}}:=1\otimes\mathrm{d}_{\mathrm{CE}}$ where $\mathrm{d}$ is the de Rham differential on $\Omega^{k}(M)$ and $\mathrm{d}_{CE}$ is the Chevalley-Eilenberg differential on $\wedge^{n-k}\mathfrak{g}^{*}$. Introduce a $k$-form taking a value in $\wedge^{n-k}\mathfrak{g}^{*}$, ${\widehat{\mu}}_{k}\in\Omega^{k}(M,\wedge^{n-k}\mathfrak{g}^{*})$, where $k=0,\ldots,n-1$, and their formal sum, ${\widehat{\mu}}=\sum_{k=0}^{n-1}{\widehat{\mu}}_{k}$.
Definition 4.2
A homotopy momentum map ${\widehat{\mu}}$ is defined by Sign factors of the equation are different from [17]. Sign factors are arranged to original ones by proper redefinitions.
$(\mathrm{d}+\mathrm{d}_{CE}){\widehat{\mu}}=\sum_{k=0}^{n}(-1)^{n-k+1}\iota_{\rho}^{n+1-k}{\omega}.$ (23)
A momentum map on an $n$-plectic manifold [10, 19] is a homotopy momentum map. Expanded by the form degree, Equation (23) becomes the following $n$ equations,
$\mathrm{d}{\widehat{\mu}}_{n-1}=-\iota_{\rho}^{1}{\omega},$ (24)
$\mathrm{d}{\widehat{\mu}}_{n-2}+\mathrm{d}_{CE}{\widehat{\mu}}_{n-1}=\iota_{\rho}^{2}{\omega},$ (25)
@@
$\mathrm{d}{\widehat{\mu}}_{k-1}+\mathrm{d}_{CE}{\widehat{\mu}}_{k}=(-1)^{n-k+1}\iota_{\rho}^{n+1-k}{\omega},$ (26)
@@
$\mathrm{d}_{CE}{\widehat{\mu}}_{0}=(-1)^{n+1}\iota_{\rho}^{n+1}{\omega}.$ (27)
The Lie group action induces an action Lie algebroid structure on the trivial bundle $E=M\times\mathfrak{g}$. The Lie algebroid differential on the action Lie algebroid is ${}^{E}\mathrm{d}=\mathrm{ad}^{*}_{\rho}+\mathrm{d}_{CE}$, where $\mathrm{ad}^{*}$ is the following action on $\alpha$ induced from the Lie algebra action on $M$,
$\mathrm{ad}^{*}_{\rho}\alpha(e_{1},\ldots,e_{m+1})$ ::=:= $\sum_{i=1}^{m+1}(-1)^{i-1}\rho(e_{i})\alpha(e_{1},\ldots,\check{e_{i}},\ldots,e_{m+1}),$ (28)
for $e_{i}\in\mathfrak{g}$, We can rewrite $n$ equations (24)–(27) to equivalent $n$ equations [22],
$\mathrm{d}\mu_{n-1}=-\iota_{\rho}^{1}{\omega},$ (29)
$\mathrm{d}\mu_{n-2}+{}^{E}\mathrm{d}\mu_{n-1}=-\iota_{\rho}^{2}{\omega},$ (30)
@@
$\mathrm{d}\mu_{k-1}+{}^{E}\mathrm{d}\mu_{k}=-\iota_{\rho}^{n+1-k}{\omega},$ (31)
@@
${}^{E}\mathrm{d}\mu_{0}=-\iota_{\rho}^{n+1}{\omega},$ (32)
where ${\widehat{\mu}}_{k}=(-1)^{n-k+1}\mu_{k}$. The last equation (32) shows that $J=\mu_{0}$ is a compatible $E$-$n$-form.
Proposition 4.3
@@
Example 4.6 (Momentum sections)
@@
Suppose that a base manifold $M$ is a pre-symplectic manifold with a pre-symplectic form $\omega$. Moreover, suppose a Lie algebroid $(E,\rho,[-,-])$ over $M$. In order to define a momentum section, we introduce a connection on the vector bundle $E$, $\nabla:\Gamma(E)\rightarrow\Gamma(T^{*}M\otimes E)$, which is a ${\mathbb{R}}$-linear map satisfying the Leibniz rule,
$\nabla(fe)=f\nabla e+(\mathrm{d}f)\otimes e,$ (33)
@@
$\mathrm{d}\langle\mu,\,e\rangle=\langle\nabla\mu,\,e\rangle+\langle\mu,\,\nabla e\rangle,$ (34)
@@
Definition 4.4
A section $\mu\in\Gamma(E^{*})$ of $E^{*}$ is called a bracket-compatible momentum section if $\mu$ satisfies the following two conditions. §§§One more condition $\nabla^{2}\mu=0$ is imposed in [5]. The condition is dropped in this paper, since it is not necessary in our context.
$\nabla\mu=-\iota_{\rho}\omega,$ (35)
${}^{E}\mathrm{d}\mu=-\iota_{\rho}^{2}{\omega}.$ (36)
The second equation (36) is nothing but the condition (11) of a compatible $E$-$1$-form for $J=\mu$.
Proposition 4.5
@@
Example 4.7 (Homotopy momentum sections)
@@
Definition 4.6
Let $E$ be a Lie algebroid over a smooth manifold $M$ and $E^{\prime}$ be a vector bundle over the same base manifold $M$. An $E$-connection on a vector bundle $E^{\prime}$ with respect to the Lie algebroid $E$ is a ${\mathbb{R}}$-linear map ${}^{E}\nabla:\Gamma(E^{\prime})\rightarrow\Gamma(E^{*}\otimes E^{\prime})$ satisfying
${}^{E}\nabla_{e}(fe^{\prime})=f{}^{E}\nabla_{e}e^{\prime}+(\rho(e)f)e^{\prime},$ (37)
@@
The standard $E$-connection on $E^{\prime}=E$, ${}^{E}\nabla:\Gamma(E)\rightarrow\Gamma(E^{*}\otimes E)$ is defined by
${}^{E}\nabla_{e}e^{\prime}:=\nabla_{\rho(e)}{e^{\prime}}+[e,e^{\prime}],$ (38)
@@
${}^{E}\nabla_{e}e^{\prime}=\nabla_{\rho(e)}{e^{\prime}}+[e,e^{\prime}]-\chi(e,e^{\prime}),$ (39)
@@
If a normal connection $\nabla$ on $E$ as a vector bundle is given, we can define the following canonical $E$-connection for the tangent bundle $E^{\prime}=TM$. An (canonical) $E$-connection on a tangent bundle, ${}^{E}\nabla:\Gamma(TM)\rightarrow\Gamma(E^{*}\otimes TM)$ is defined by
${}^{E}\nabla_{e}v$ ::=:= ${\mathcal{L}}_{\rho(e)}v+\rho(\nabla_{v}e)=[\rho(e),v]+\rho(\nabla_{v}e),$ (40)
@@
Given a normal connection $\nabla$, a covariant derivative is generalized to the derivation on the space of differential forms, $\mathrm{d}^{\nabla}:\Omega^{k}(M,\wedge^{m}E^{*})\rightarrow\Omega^{k+1}(M,\wedge^{m}E^{*})$, called an exterior covariant derivativeThe exterior covariant derivative is also denoted by $\nabla$ in [22].. Similarly, given an $E$-connection ${}^{E}\nabla$, the Lie algebroid differential ${}^{E}\mathrm{d}$ can be generalized to the operation on the space $\Gamma(\wedge^{k}E^{\prime}\otimes\wedge^{m}E^{*})$, ${}^{E}\mathrm{d}^{\nabla}:\Gamma(\wedge^{k}E^{\prime}\otimes\wedge^{m}E^{*})\rightarrow\Gamma(\wedge^{k}E^{\prime}\otimes\wedge^{m+1}E^{*})$. For instance, see [1]. The concrete equation for $E^{\prime}=T^{*}M$ is as follows.
Definition 4.7
For $\Omega^{k}(M,\wedge^{m}E^{*})=\Gamma(\wedge^{k}T^{*}M\otimes\wedge^{m}E^{*})$, the $E$-exterior covariant derivative ${}^{E}\mathrm{d}^{\nabla}:\Omega^{k}(M,\wedge^{m}E^{*})\rightarrow\Omega^{k}(M,\wedge^{m+1}E^{*})$ is defined by
${}^{E}\mathrm{d}^{\nabla}\alpha(e_{1},\ldots,e_{m+1})$ ::=:= $\sum_{i=1}^{m+1}(-1)^{i-1}{}^{E}\nabla_{e_{i}}(\alpha(e_{1},\ldots,\check{e_{i}},\ldots,e_{m+1}))$ (41)
$+\sum_{1\leq i<j\leq m+1}(-1)^{i+j}\alpha([e_{i},e_{j}],e_{1},\ldots,\check{e_{i}},\ldots,\check{e_{j}},\ldots,e_{m+1}),$
@@
Let $\mu_{k}\in\Omega^{k}(M,\wedge^{n-k}E^{*})$ be a $k$-form taking values in $\wedge^{n-k}E^{*}$, where $k=0,\ldots,n-1$. Both the exterior covariant derivative $\nabla$ and the $E$-exterior covariant derivative ${}^{E}\mathrm{d}^{\nabla}$ act on the space $\Omega^{k}(M,\wedge^{n-k}E^{*})$.
Definition 4.8

### v4

@@ 3.2 Gauge transformations
Recall that $\iota^{n+1}_{\rho}\omega$ is closed with respect to the Lie algebroid differential ${{}^{E}\mathrm{d}}$, ${{}^{E}\mathrm{d}}\bigl{(}\iota^{n+1}_{\rho}\omega\bigr{)}=0$ from consistency with equations (3.1) and (3.2). Obvious equivalence relation of compatible $E$-$n$-forms is difference of ${}^{E}\mathrm{d}$-exact terms.
Definition 3.4.
For a Lie algebroid $E$, two compatible $E$-$n$-forms $J$ and $J^{\prime}$ are equivalent if there exists an $E$-$(n-1)$-form $K\in\Gamma\bigl{(}\wedge^{n-1}E^{*}\bigr{)}$ satisfying $J^{\prime}=J+{}^{E}\mathrm{d}K$.
@@
${{}^{E}\mathrm{d}}\bigl{(}\iota_{\rho}^{n}\lambda\bigr{)}=\iota_{\rho}^{n+1}\mathrm{d}\lambda$
@@
${{}^{E}\mathrm{d}}J^{\prime}=-\iota_{\rho}^{n+1}\omega^{\prime}=-\iota_{\rho}^{n+1}\omega-(\iota_{\rho})^{n+1}\mathrm{d}\lambda=-\iota_{\rho}^{n+1}\omega-{{}^{E}\mathrm{d}}\bigl{(}\iota_{\rho}^{n}\lambda\bigr{)}={{}^{E}\mathrm{d}}\bigl{(}J-\iota_{\rho}^{n}\lambda\bigr{)}.$
@@
In Section 4, we discuss gauge equivalence of the twisted Poisson structure in Example 4.2.
@@ 4 Examples
The compatible $E$-$n$-form (3.1) is considered as a universal compatibility condition of a Lie algebroid structure with a pre-multisymplectic form. It appears in various geometries. In this section, we list up examples.
Example 4.1 (Liouville $1$-forms).
Take $n=1$ and let $(M,\omega)$ be a symplectic manifold and $E=TM$ be a tangent bundle. $TM$ is a tangent Lie algebroid with the anchor map $\rho=\mathrm{id}$ (see Example 2.3). An $E$-$1$-form $J$ is a normal differential form on $M$, and a Lie algebroid differential ${}^{E}\mathrm{d}=\mathrm{d}_{\mathrm{dR}}$ is the de Rham differential. The compatible condition (3.1) is
$\mathrm{d}_{\mathrm{dR}}J=-\omega,$
@@
Example 4.2 (Poisson and twisted Poisson structures).
Let $(\pi,H)$ be a twisted Poisson structure on $M$ (see Example 2.6). Note that it reduces to a Poisson structure if $H=0$. The cotangent bundle $T^{*}M$ has a Lie algebroid structure as explained in Examples 2.5 and 2.6. Formulas of the twisted Poisson structures and induced Lie algebroids are in [32, 40]. Equation (2.3) is equivalent to ${}^{E}\mathrm{d}\pi=-\bigl{\langle}\otimes^{3}\pi,H\bigr{\rangle}$, where ${}^{E}\mathrm{d}$ is the Lie algebroid differential of this Lie algebroid $T^{*}M$. Thus, $J=\pi$ is a compatible $TM$-$2$-form on a pre-$2$-plectic manifold with a pre-$2$-plectic form $H$.
Next let us analyze gauge equivalence. A gauge transformation of the twisted Poisson structure is $H^{\prime}=H+\mathrm{d}B$ with a $2$-form $B$. Then, the compatible $E$-$2$-form with respect to $H^{\prime}$ is $\pi^{\prime}=\pi-\bigl{\langle}\otimes^{2}\pi,B\bigr{\rangle}$ from Definition 3.5.
@@
In the standard definition, equation (2.3) gives the Lie algebroid structure on $T^{*}M$ and a compatible $E$-$n$-form simultaneously. If the gauge transformation $\pi^{\prime}=\pi(1+\pi B)^{-1}$ consistently deforms the Lie algebroid structure defined by $(\pi^{\prime},H)$ and a compatible $E$-$n$-form to ${J^{\prime}=\pi^{\prime}}$ simultaneously.
However, our gauge transformation $\pi^{\prime}=\pi-\bigl{\langle}\otimes^{2}\pi,B\bigr{\rangle}$ only deforms the compatible $E$-$n$-form to $J^{\prime}=\pi^{\prime}$ under the fixed Lie algebroid constructed from $(\pi,H)$. The reason is for generalizations to higher pre-$n$-plectic cases. In a pre-$n$-plectic case, the simultaneous deformation of the Lie algebroid and the compatible $E$-$n$-form like the twisted Poisson structure does not work. The twisted Poisson structure is a very special case and our gauge transformation is applicable to general $E$-$n$-form on a higher $n$-plectic manifold.
Example 4.3.
A twisted $R$-Poisson structure $(\pi,H,J)$ on a smooth manifold $M$ is defined as follows [13]. [^1] Since $J$ is denoted by $R$ in the paper [13], it is called the $R$-Poisson structure. $M$ is a smooth manifold and $\pi\in\Gamma\bigl{(}\wedge^{2}TM\bigr{)}$ is a Poisson bivector field. $H$ is a closed $(n+1)$-form, and $J\in\Gamma(\wedge^{n}TM)$ is an $n$-multivector field. As explained in Example 2.5, the Poisson bivector field $\pi$ induces a Lie algebroid structure on $T^{*}M$, where $\rho=-\pi^{\sharp}$ and the Lie bracket is given by (2.2). Under this Lie algebroid structure, $(H,J)$ is the twisted $R$-Poisson structure if it satisfies
${}^{E}\mathrm{d}J=(-1)^{n}\bigl{\langle}\otimes^{n+1}\pi,H\bigr{\rangle}.$
@@
Example 4.4 (momentum maps).
We assume a smooth action of a Lie group $G$ on a symplectic manifold $(M,\omega)$. Let $\mathfrak{g}$ be a Lie algebra for $G$. The Lie group action induces the Lie algebra action $\rho\colon\mathfrak{g}\rightarrow TM$. A map $\mu\colon M\rightarrow\mathfrak{g}^{*}$ is called a momentum map if it satisfies
$\mathrm{d}\mu(e)=-\iota_{\rho(e)}\omega,$ (4.1)
$\rho(e_{1})\langle\mu,\,e_{2}\rangle=\langle\mu,\,[e_{1},e_{2}]\rangle$ (4.2)
for $e,e_{1},e_{2}\in\mathfrak{g}$. Here $\langle$-,\,$\rangle$ - , - ⟩ is the pairing of $\mathfrak{g}^{*}$ and $\mathfrak{g}$. As explained in Example 2.4, the Lie group action induces an action Lie algebroid structure on a trivial bundle $E=M\times\mathfrak{g}$. The anchor map is induced from the map of the Lie algebra action as $\rho\colon M\times\mathfrak{g}\rightarrow TM$. The momentum map is regarded as a section of the dual trivial bundle $\mu\in\Gamma(M\times\mathfrak{g}^{*})$. Under the condition (4.1), equation (4.2) is changed to [6]
${}^{E}\mathrm{d}\mu(e_{1},e_{2})=-\iota_{\rho}^{2}\omega(e_{1},e_{2}),$
@@
Proposition 4.5.
A momentum map $\mu\colon M\rightarrow\mathfrak{g}^{*}$ is a compatible $E$-$1$-form for the action Lie algebroid induced from a Lie group action on $M$.
Example 4.6 (homotopy moment(um) maps).
[10] Let $(M,{\omega})$ be an $n$-plectic manifold. Assume an action of a Lie group $G$ on $M$ as in the case of the momentum map. The action of $G$ induces the corresponding infinitesimal Lie algebra action on $M$ as vector fields, $\rho\colon\mathfrak{g}\rightarrow TM$. We consider the space, $\Omega^{k}\bigl{(}M,\wedge^{n-k}\mathfrak{g}^{*}\bigr{)}=\Omega^{k}(M)\otimes\wedge^{n-k}\mathfrak{g}^{*}$, and two differentials ${\mathrm{d}:=\mathrm{d}\otimes 1}$ and ${\mathrm{d}_{\mathrm{CE}}:=1\otimes\mathrm{d}_{\mathrm{CE}}}$ where $\mathrm{d}$ is the de Rham differential on $\Omega^{k}(M)$ and $\mathrm{d}_{CE}$ is the Chevalley–Eilenberg differential on $\wedge^{n-k}\mathfrak{g}^{*}$. Introduce a $k$-form taking a value in $\wedge^{n-k}\mathfrak{g}^{*}$, ${\widehat{\mu}}_{k}\in\Omega^{k}\bigl{(}M,\wedge^{n-k}\mathfrak{g}^{*}\bigr{)}$, where $k=0,\dots,n-1$, and their formal sum, ${\widehat{\mu}}=\sum_{k=0}^{n-1}{\widehat{\mu}}_{k}$.
Definition 4.7.
A homotopy momentum map ${\widehat{\mu}}$ is defined by [^2] Sign factors of the equation are different from [10]. Sign factors are arranged to original ones by proper redefinitions.
$(\mathrm{d}+\mathrm{d}_{CE}){\widehat{\mu}}=\sum_{k=0}^{n}(-1)^{n-k+1}\iota_{\rho}^{n+1-k}{\omega}.$ (4.3)
A momentum map on an $n$-plectic manifold [11, 19] is a homotopy momentum map. Expanded by the form degree, equation (4.3) becomes the following $n$ equations,
$\mathrm{d}{\widehat{\mu}}_{n-1}=-\iota_{\rho}^{1}{\omega},$
$\mathrm{d}{\widehat{\mu}}_{n-2}+\mathrm{d}_{CE}{\widehat{\mu}}_{n-1}=\iota_{\rho}^{2}{\omega},$
@@
$\mathrm{d}{\widehat{\mu}}_{k-1}+\mathrm{d}_{CE}{\widehat{\mu}}_{k}=(-1)^{n-k+1}\iota_{\rho}^{n+1-k}{\omega},$
@@
$\mathrm{d}_{CE}{\widehat{\mu}}_{0}=(-1)^{n+1}\iota_{\rho}^{n+1}{\omega}.$ (4.4)
The Lie group action induces an action Lie algebroid structure on the trivial bundle ${E\!=\!M\!\times\!\mathfrak{g}}$. The Lie algebroid differential on the action Lie algebroid is ${}^{E}\mathrm{d}=\operatorname{ad}^{*}_{\rho}+\mathrm{d}_{CE}$, where $\operatorname{ad}^{*}$ is the following action on $\alpha$ induced from the Lie algebra action on $M$,
$\operatorname{ad}^{*}_{\rho}\alpha(e_{1},\dots,e_{m+1}):=\sum_{i=1}^{m+1}(-1)^{i-1}\rho(e_{i})\alpha(e_{1},\dots,\check{e_{i}},\dots,e_{m+1})$
for $e_{i}\in\mathfrak{g}$, We can rewrite $n$ equations (4.4) to equivalent $n$ equations [22],
$\mathrm{d}\mu_{n-1}=-\iota_{\rho}^{1}{\omega},$
$\mathrm{d}\mu_{n-2}+{}^{E}\mathrm{d}\mu_{n-1}=-\iota_{\rho}^{2}{\omega},$
@@
$\mathrm{d}\mu_{k-1}+{}^{E}\mathrm{d}\mu_{k}=-\iota_{\rho}^{n+1-k}{\omega},$
@@
${}^{E}\mathrm{d}\mu_{0}=-\iota_{\rho}^{n+1}{\omega},$ (4.5)
where ${\widehat{\mu}}_{k}=(-1)^{n-k+1}\mu_{k}$. The last equation (4.5) shows that $J=\mu_{0}$ is a compatible $E$-$n$-form.
Proposition 4.8.
@@
Example 4.9 (momentum sections).
@@
Suppose that a base manifold $M$ is a pre-symplectic manifold with a pre-symplectic form $\omega$. Moreover, suppose a Lie algebroid $(E,\rho,[-,-])$ over $M$. In order to define a momentum section, we introduce a connection on the vector bundle $E$, $\nabla\colon\Gamma(E)\rightarrow\Gamma(T^{*}M\otimes E)$, which is a ${\mathbb{R}}$-linear map satisfying the Leibniz rule,
$\nabla(fe)=f\nabla e+(\mathrm{d}f)\otimes e$
@@
$\mathrm{d}\langle\mu,\,e\rangle=\langle\nabla\mu,\,e\rangle+\langle\mu,\,\nabla e\rangle$
@@
Definition 4.10.
A section $\mu\in\Gamma(E^{*})$ of $E^{*}$ is called a bracket-compatible momentum section if $\mu$ satisfies the following two conditions: [^3] One more condition $\nabla^{2}\mu=0$ is imposed in [6]. The condition is dropped in this paper, since it is not necessary in our context.
$\nabla\mu=-\iota_{\rho}\omega,\qquad{}^{E}\mathrm{d}\mu=-\iota_{\rho}^{2}{\omega}.$ (4.6)
The second equation (4.6) is nothing but the condition (3.1) of a compatible $E$-$1$-form for ${J=\mu}$.
Proposition 4.11.
@@
Example 4.12 (homotopy momentum sections).
@@
Definition 4.13.
Let $E$ be a Lie algebroid over a smooth manifold $M$ and $E^{\prime}$ be a vector bundle over the same base manifold $M$. An $E$-connection on a vector bundle $E^{\prime}$ with respect to the Lie algebroid $E$ is a ${\mathbb{R}}$-linear map ${}^{E}\nabla\colon\Gamma(E^{\prime})\rightarrow\Gamma(E^{*}\otimes E^{\prime})$ satisfying
${}^{E}\nabla_{e}(fe^{\prime})=f{}^{E}\nabla_{e}e^{\prime}+(\rho(e)f)e^{\prime}$
@@
The standard $E$-connection on $E^{\prime}=E$, ${}^{E}\nabla\colon\Gamma(E)\rightarrow\Gamma(E^{*}\otimes E)$ is defined by
${}^{E}\nabla_{e}e^{\prime}:=\nabla_{\rho(e)}{e^{\prime}}+[e,e^{\prime}]$
@@
${}^{E}\nabla_{e}e^{\prime}=\nabla_{\rho(e)}{e^{\prime}}+[e,e^{\prime}]-\chi(e,e^{\prime})$
@@
If a normal connection $\nabla$ on $E$ as a vector bundle is given, we can define the following canonical $E$-connection for the tangent bundle $E^{\prime}=TM$. An (canonical) $E$-connection on a tangent bundle, ${}^{E}\nabla\colon\Gamma(TM)\rightarrow\Gamma(E^{*}\otimes TM)$ is defined by
${}^{E}\nabla_{e}v:={\mathcal{L}}_{\rho(e)}v+\rho(\nabla_{v}e)=[\rho(e),v]+\rho(\nabla_{v}e),$
@@
Given a normal connection $\nabla$, a covariant derivative is generalized to the derivation on the space of differential forms, $\mathrm{d}^{\nabla}\colon\Omega^{k}(M,\wedge^{m}E^{*})\rightarrow\Omega^{k+1}(M,\wedge^{m}E^{*})$, called an exterior covariant derivative. [^4] The exterior covariant derivative is also denoted by $\nabla$ in [22]. Similarly, given an $E$-connection ${}^{E}\nabla$, the Lie algebroid differential ${}^{E}\mathrm{d}$ can be generalized to the operation on the space
$\Gamma\bigl{(}\wedge^{k}E^{\prime}\otimes\wedge^{m}E^{*}\bigr{)},\qquad{}^{E}\mathrm{d}^{\nabla}\colon\Gamma\bigl{(}\wedge^{k}E^{\prime}\otimes\wedge^{m}E^{*}\bigr{)}\rightarrow\Gamma\bigl{(}\wedge^{k}E^{\prime}\otimes\wedge^{m+1}E^{*}\bigr{)}.$
For instance, see [1]. The concrete equation for $E^{\prime}=T^{*}M$ is as follows.
Definition 4.14.
For
$\Omega^{k}(M,\wedge^{m}E^{*})=\Gamma\bigl{(}\wedge^{k}T^{*}M\otimes\wedge^{m}E^{*}\bigr{)},$
the $E$-exterior covariant derivative ${}^{E}\mathrm{d}^{\nabla}\colon\Omega^{k}(M,\wedge^{m}E^{*})\rightarrow\Omega^{k}\bigl{(}M,\wedge^{m+1}E^{*}\bigr{)}$ is defined by
${}^{E}\mathrm{d}^{\nabla}\alpha(e_{1},\dots,e_{m+1}):={}$ $\sum_{i=1}^{m+1}(-1)^{i-1}{}^{E}\nabla_{e_{i}}(\alpha(e_{1},\dots,\check{e_{i}},\dots,e_{m+1}))$
$+\sum_{1\leq i<j\leq m+1}(-1)^{i+j}\alpha([e_{i},e_{j}],e_{1},\dots,\check{e_{i}},\dots,\check{e_{j}},\dots,e_{m+1})$
@@
Let $\mu_{k}\in\Omega^{k}\bigl{(}M,\wedge^{n-k}E^{*}\bigr{)}$ be a $k$-form taking values in $\wedge^{n-k}E^{*}$, where $k=0,\dots,n-1$. Both the exterior covariant derivative $\nabla$ and the $E$-exterior covariant derivative ${}^{E}\mathrm{d}^{\nabla}$ act on the space ${\Omega^{k}\bigl{(}M,\wedge^{n-k}E^{*}\bigr{)}}$.
Definition 4.15.